OpenStudy (anonymous):

please solve: factoring by groupings :3x+6+xy+2y,,-5x+1+y+5xy,,6x+4+3xy+2y,,x2+2x+xy+2y,,7x+14+xy2+2y2

OpenStudy (anonymous):

3(x+2)+y(x+2)y=(x+2)(3+y)

OpenStudy (anonymous):

\[3x + 6 + xy + 2y = 3(x+2) + y(x + 2) = (x+2)(y+3)\]

OpenStudy (anonymous):

6x+4+3xy+2y = 3xy+2y+6x+4 = (3x+2)(y+2)